OpenStudy (anonymous):

Check my answers? Factor this polynomial using the GFC: 4x + 2 A 4(x + 1) B 2(4x + 2) C 2(2x + 1) D 2(2x + 2) I answered b Factor this polynomial using the GFC: 21x - 14 A x(21 - 14) B 7(3x - 1) C 21x(1 - 14) D 7(3x - 2) I answered c Factor this polynomial using the GFC: 12x^2 - 3x + 6 A 3(4x^2 - 3x + 2) B 3(4x^2 - x + 2) C 4(3x^2 - x + 1) D 3(4x^2 + 1x + 2) I answered a Factor this polynomial using the GFC: -4x^2 - 16x - 2 A -4(x^2 + 4x - 1) B -2(2x^2 + 8x + 1) C -2(2x^2 - 8x +1) D -2(2x^2 - 8x - 1) I answered c

OpenStudy (zzr0ck3r):

what factor dies 4x and 2 have in common?

OpenStudy (zzr0ck3r):

what factor does*

OpenStudy (luigi0210):

you might wanna check your answers again ^_^'

OpenStudy (zzr0ck3r):

2(4x+2) = 2*4x+2*4 = 8x+4 != 4x+2

OpenStudy (anonymous):

i would double check your answers:)

OpenStudy (anonymous):

You two really aren't much help if you don't tell me what I did wrong or what I got wrong.. Zzr0ck3r which question are you talking about? I'm kinda lost